# Multiple Instance Verification

### v1

@@ 6.3 Results for fact extraction and verification
{NiceTabular}
@l@  c@ c@ c@ c@ c@ AUROCAccuracyPrecisionRecallF1-score
Baseline0.864$\pm$0.0040.778$\pm$0.0030.780$\pm$0.0000.777$\pm$0.0030.777$\pm$0.003
GABMIL0.864$\pm$0.0030.777$\pm$0.0020.780$\pm$0.0000.777$\pm$0.0030.777$\pm$0.003
PMA0.855$\pm$0.0020.770$\pm$0.0020.773$\pm$0.0030.770$\pm$0.0000.770$\pm$0.000
MSA0.857$\pm$0.0010.771$\pm$0.0030.767$\pm$0.0030.767$\pm$0.0030.767$\pm$0.003
MI-Net 0.857$\pm$0.002 0.769$\pm$0.002 0.770$\pm$0.000 0.770$\pm$0.000 0.770$\pm$0.000
Bi-LSTM 0.860$\pm$0.002 0.771$\pm$0.001 0.770$\pm$0.000 0.770$\pm$0.000 0.770$\pm$0.000
CAP-VEMA(ours)0.886$\pm$0.0020.807$\pm$0.0030.813$\pm$0.0030.807$\pm$0.0030.807$\pm$0.003
@@
Table 8: Fact extraction and verification (FEVER): evaluating the explanations (mean and standard errors). The best performance is shown in bold.
{NiceTabular}
@l@cc@ Avg. i-AUROC Avg. i-AP
@@
[…]
Performance on the FEVER task is shown in Tables 6.3 and 8. As in the other tasks, CAP-based models outperform the baseline and benchmarks, with statistical significance and by sizable margins—particularly large margins in the quality of explanations. Interestingly, DBA provides slight but consistent improvements over VEMA across all measures. Similar to the second task, there is some inconsistency between the quality of classifications and explanations when comparing the benchmarks to the baseline.
@@ 6.4 Ablation study
{NiceTabular}
[…]
@@
{NiceTabular}
[…]
@@
(c) Incremental effects of model components by progressively building CAP-VEMA and CAP-DBA from the baseline, in the QMNIST task.
{NiceTabular}
[…]
@@
[…]
(d) Building CAP-VEMA from the baseline.
{NiceTabular}
[…]
@@
[…]
(e) Building CAP-DBA from the baseline.
(f) Incremental effects of model components by progressively building CAP-VEMA and CAP-DBA from the baseline, in the signature verification task.
{NiceTabular}
[…]
@@
[…]
(g) Building CAP-VEMA from the baseline.
{NiceTabular}
[…]
@@
[…]
(h) Building CAP-DBA from the baseline.
@@
(i) Incremental effects of model components by progressively building CAP-VEMA and CAP-DBA from the baseline, in the FEVER task.
@@
To understand the effects of components from our models, it is instructive to consider the incremental effects of progressively adding CAP components to the baseline, one component at a time, using both VEMA and DBA attention functions. The results of the incremental effects on performance are presented in Tables 6.4, 6.4, 6.4, for the three tasks respectively.
@@
The QMNIST task is synthesized by generating each exemplar using the same digit, making instances within a target bag all look alike. As such, it may be a well-controlled test of the effectiveness of our approach. Table 6.4 demonstrates that the new functional forms of attention, VEMA and DBA, contribute far more to quality of explanations than the other components, and a large portion of the classification enhancement. Note that in this task, the baseline’s performance already significantly exceeds that of the three benchmarks (cf. Table 6.1.1 and 3), possibly due to its attention incorporating the query instance to some extent. Yet on top of the baseline’s performance, VEMA/DBA continue to bring substantial improvement. This strongly supports the effectiveness of our proposed attention functions to tackle the challenges brought by this task. Particularly for DBA, despite its simpler form, we note that its performance is comparable to that of VEMA.
@@
For the two other tasks, signature verification and FEVER, the overall patterns are similar: the new attention functions in conjunction with “squeeze and co-excitation” (SCE) contribute the most to model performance, for both classification and quality of explanations. Again, the functional form of attention contributes the most by far. They and SCE are typically the top two contributors to classification performance.
@@
Results demonstrate that the majority of uplift in capability to identify key instances can be attributed to the new functional forms of attention, VEMA and DBA. In conjunction with SCE, they also contribute the most to classification enhancement.
@@
On the other hand, the contribution of the multi-head projection is mixed. When moving from the “non-headed” version (without linear projections) to a multi-head version, the performance differences vary across tasks and measures. Additional analyses provided in Appendix D isolate effects from the multi-head mechanism by comparing models with and without multiple heads. Those analyses demonstrate the limited contribution from the multi-head mechanism, consistent with the results in this section. The reason may be that, in these tasks, each instance is typically an independent and complete object with a holistic—not partial—feature representation, while the multi-head mechanism is designed “to jointly attend to information from different representation subspaces at different positions” [34]. Therefore, given the presence of other model components, multi-headedness alone appears to add little value in the tasks considered in this paper. [^10] Note that the transformer used separate multi-head projections for each “query, key, value” triplet in attention. Based on the ablation results, we chose not to use separate projections for our experiments. Instead, we opted for less model complexity, by employing a single, shared linear projection in the multi-head formulation of CAP.
@@
Finally, from Tables 6.4 and 6.4 we observe that there are consistent improvements brought by the component of multi-head “pre-aggregation LayerNorm” (cf. Equations (8a), (8b), as the last element in the CAP architecture), compared to no normalization. To check the robustness of CAP to such normalization, we tested the sensitivity of our model results to an alternative form of LayerNorm, shown in Appendix E. This alternative is a standard LayerNorm without multi-headness, akin to the post-attention LayerNorm adopted in the transformer model, applied after bag aggregation and after all heads are concatenated in Equations (8a) and (8b). Henceforth, we call it “post-aggregation LayerNorm”. The results in Appendix E show similar performance of CAP-based models when the pre-aggregation LayerNorm is replaced with the post-aggregation LayerNorm, and thus demonstrate the robustness of CAP to different forms of output normalization.
@@ 7 Conclusions and future research
We introduce MI verification, which combines verification and MIL, and show theoretically and empirically that a lack of information about the query instance in attention is undesirable when pooling the target bag into a bag-level representation. To empirically demonstrate this, we construct a simple baseline model and three benchmark models, adapted from SOTA attention-based MIL methods, that omit the query in attention, and create a new approach named “Cross Attention Pooling” (CAP) that explicitly considers the query in attention, along with two new attention functions within the CAP framework. Results on three different tasks show that the benchmark methods are not better, and sometimes significantly worse, than the simple baseline. In contrast, the CAP-based methods outperform the baseline and benchmarks by sizeable margins in terms of both classification and quality of explanations. Ablation studies confirm the superior ability of the new attention functions to identify key instances and establish the contributions of the key components of the CAP architecture.
@@
The abstract representation of “MI verification”, as distinct from traditional MIL opens up the potential for research on a broad range of problems and applications. For example, either the “query” or the “target” may be a single instance, bag of instances, or bag of bags, and they may exhibit different modalities—for example, one may be textual and the other may consist of images. We believe our work provides a generic framework that may inspire research in directions such as new architectures, attention mechanisms, real-world applications, and methods for generating explanations and evaluating their quality.
@@
As a first step towards tackling MI verification, our paper also offers future opportunities connecting to some domain-specific, well-established research areas. For example, while this paper considers CV datasets that enable well-controlled experiments and disentangling factors without compromising on the difficulty of tasks, applications of MI verification on more complex datasets such as ImageNet or MS-COCO may lead to research of weakly-supervised one-shot object detection or weakly-supervised cross-domain retrieval (cf. Section 3). For the NLP domain, a comparison of our approach to other methods specifically designed for particular tasks, e.g., FEVER, may also be an interesting topic for prospective research (cf. Section 2). Interpretability of weakly supervised learning and its evaluation is yet another avenue that this work can be extended to. Those research topics are sufficiently significant and challenging to warrant dedicated, standalone future studies.
[…]
@@ Appendix A Proof of Proposition 1
Proof.
[Proposition 1: Variables incorporated in the attention scores]
@@
This concludes the proof of Inequality (4). By symmetry, Inequality (5) can also be shown in a similar manner, which completes the proof of Proposition 1. ∎
@@ Appendix B Details of data and training process
Prior to full training of all network parameters, we adopt an initial learning stage where we freeze the feature extractor’s model weights and train only the parameters specific to the baseline, benchmarks. This initial phase is stopped when the accuracy on the validation dataset does not improve for two epochs.
Subsequently, when all parameters are trained end-to-end, we also adopt early stopping based on validation accuracy, with different stopping criteria for different tasks as discussed below. The optimization for training is conducted using the RMSprop optimizer, with the same parameters for all models: $\rho=0.9,\epsilon=$1e-7. Due to a well-known behavior of “batch normalization” (BatchNorm) causing degradation in performance during transfer learning [19], we freeze the BatchNorm layers, if any, in a feature extractor, typically for CV tasks. More precisely, the BatchNorm moving average and the trainable parameters are not updated during training. [^11] We also tried allowing parameters to be trainable but not updating the moving average parameters and found similar experimental results.
For each task, we conducted three rounds of experiments for all models considered in this paper. All experiments were run on a cluster of four NVIDIA RTX A6000 GPUs, of which the duration varies depending on early stopping triggers. Roughly speaking, a single round of training of one model takes approximately one hour for the QMNIST task, two to three hours for the FEVER task, and three to five hours for the signature verification task.
All models were developed using Tensorflow 2.9.3, with some use of “TensorFlow official models” 2.9.0 [40] (Apache License 2.0) and scikit-learn 1.2.0 [27] (BSD License). The original license/terms-of-use of the assets (i.e., raw data, pre-trained feature extractors) used in our research are listed alongside each asset mentioned below.
@@ B.1 QMNIST handwriting verification
We collected QMNIST data (BSD-style license, https://github.com/facebookresearch/qmnist/blob/main/LICENSE) according to [38] and from links therein. QMNIST was created to be similar in structure to MNIST, with a training set composed of 60,000 images of digits (“Train/60K”) and a test set of 10,000 digits (“Test/10K”). These two sets turn out to have non-overlapping writer-IDs, which is ideal for our requirement that test writer’s handwriting is not seen during training. We constructed a training dataset by randomly selecting from the “Train/60K” set, including both the query and target bag. To construct the validation and test datasets, we first split “Test/10K” into two sets with non-overlapping writer-IDs and then drew random samples from each set. By construction, the sample sizes of the train/dev/test datasets are 21,509/2,408/2,253 respectively (the same sample sizes for all rounds of experiments were used to draw random samples), and the proportions of their exemplar labels are approximately 50:50. In an exemplar, the bag size $N$ varies between 3 and 25, with a mean of $6.9$ instances per target bag and a variance of $6.4$ approximately.
We employed ResNet-18 [14] as the feature extractor, initialized by a set of pre-trained weights based on ImageNet. The pre-trained weights were downloaded from https://pypi.org/project/image-classifiers (MIT License). Prior to inputting the QMNIST images into ResNet-18, we resized them from their original size of $(28\times 28)$ to $(32\times 32)$, because this is the minimum input size required by ResNet-18. After putting an image through ResNet-18, we obtained its feature vector by applying a “Global Average Pooling” (GAP) on the output of the feature extractor’s penultimate layer.
The learning rate of the RMSprop optimizer was piece-wise constant: 1e-4 for the first 5 epochs, 5e-5 for the next 15 epochs, and 2e-5 for the remaining epochs. The mini-batch size was 768, and the early-stopping criterion was non-improving validation accuracy for 30 epochs. The number of heads was two for any verification models that support multi-head attention.
@@ B.2 Signature verification against multiple anchors
We collected raw images of signatures, both authentic and forged, based on [24] and from the link http://www.iapr-tc11.org/mediawiki/index.php/ICDAR_2011_Signature_Verification_Competition_(SigComp2011) (data disclaimer: http://www.iapr-tc11.org/dataset/ICDAR_SignatureVerification/SigComp2011/disclaimer.pdf). We resized all signature images from their original sizes to $(512\times 256)$ and converted them from RGB to black-and-white. To construct the train, validation, and test datasets for each round of the experiment, we first split the raw data into three sets with non-overlapping writer-IDs and then drew random samples from each set. The sample size for training is around 82,000, and for validation/test is close to 10,000. By construction, the proportion of exemplar labels in all datasets is approximately 50:50. In an exemplar, the bag size $N$ varies between two and eight.
For the feature extractor, we employed one from the EfficientNetV2 family [31], called “EfficientNetV2B3”, initialized with ImageNet-pretrained weights. The pretrained weights were downloaded from https://www.tensorflow.org/versions/r2.9/api_docs/python/tf/keras/applications/efficientnet_v2/EfficientNetV2B3 (CC-BY-4.0 license). We resized all the signature images to an identical size of $(256\times 128)$. After putting an image through EfficientNetV2B3, we obtained its feature vector by applying global average pooling on the output from the penultimate layer.

### v2

@@ 6.3 Results for fact extraction and verification
AUROC Accuracy Precision Recall F1-score
@@
[…]
Table 9: Fact extraction and verification (FEVER): key instance detection performance (mean and standard errors). The best performance is shown in bold.
Avg. i-AUROC Avg. i-AP
@@
CAP-DBA 0.851 $\pm$ 0.004 0.743 $\pm$ 0.005
Performance on the FEVER task is shown in Tables 8 and 9. As in other tasks, CAP-based models outperform the baseline and benchmarks by sizable margins—particularly large margins in key instance detection performance. Interestingly, DBA provides slight but consistent improvements over VEMA across all measures. Similarly to the second task, there is some inconsistency between the quality of classifications and key instance detection when comparing the benchmarks with the baseline.
@@ 6.4 Ablation study
Model Components Classification Key instance detection
@@
Model Components Classification Key instance detection
@@
Table 10: Incremental effects of model components by progressively building CAP-VEMA and CAP-DBA from the baseline, in the QMNIST task.
Model Components Classification Key instance detection
@@
CAP-VEMA 0.819 0.673 0.995 0.993
(a) Building CAP-VEMA from the baseline.
Model Components Classification Key instance detection
@@
CAP-DBA 0.814 0.674 0.993 0.991
(b) Building CAP-DBA from the baseline.
Table 11: Incremental effects of model components by progressively building CAP-VEMA and CAP-DBA from the baseline, in the signature verification task.
Model Components Classification Key instance detection
@@
CAP-VEMA 0.886 0.807 0.835 0.719
(a) Building CAP-VEMA from the baseline.
Model Components Classification Key instance detection
@@
CAP-DBA 0.898 0.817 0.851 0.743
(b) Building CAP-DBA from the baseline.
@@
Table 12: Incremental effects of model components by progressively building CAP-VEMA and CAP-DBA from the baseline, in the FEVER task.
@@
To understand the effects of components from our models, it is instructive to consider the incremental effects of progressively adding CAP components to the baseline, one component at a time, using both VEMA and DBA attention functions. The results of the incremental effects on performance are presented in Tables 10, 11, 12, for the three tasks, respectively.
@@
The QMNIST task is synthesized by generating each exemplar using the same digit, making instances within a target bag all look alike. As such, it may be a well-controlled test of the effectiveness of our approach. Table 10 demonstrates that the new functional forms of attention, VEMA and DBA, contribute substantially more to key instance detection performance than the other components, and to a large portion of the improvement in classification performance. Note that in this task, the baseline’s performance already significantly exceeds that of the benchmarks (cf. Tables 2 and 4), possibly because its attention implicitly incorporates information from the query instance. On top of the baseline, VEMA and DBA bring further substantial improvement. This strongly supports the effectiveness of our proposed attention functions in this challenging task. We note that the performance of DBA, despite its simpler form, is comparable to that of VEMA.
@@
For the other two tasks, signature verification and FEVER, the overall patterns are similar: the new functional form of attention is typically the top contributor to model performance. The two components, the attention function and “co-excitation”, jointly account for the majority of performance, for both classification and key instance detection.
@@
These results demonstrate that most of the increased ability to identify key instances can be attributed to the new functional forms of attention, VEMA and DBA. In conjunction with co-excitation, they also contribute the most to classification enhancement.
@@
On the other hand, the contribution of the multi-head projection is mixed. When moving from the “non-headed” version (i.e., without linear projections) to a multi-head version, the performance differences vary across tasks and measures. Additional analyses provided in Appendix D isolate effects from the multi-head mechanism by comparing models with and without multiple heads. Those analyses demonstrate the limited contribution from the multi-head mechanism, consistent with the results in this section. The reason may be that, in these tasks, each instance is typically an independent and complete object with a holistic—not partial—feature representation, while the multi-head mechanism is designed “to jointly attend to information from different representation subspaces at different positions” (Vaswani et al. (2017)). Therefore, given the presence of other model components, multi-headedness alone appears to add little value in the tasks considered in our study. [^14] Note that the transformer used separate multi-head projections for each “query, key, value” triplet in attention. Based on the ablation results, we chose not to use separate projections for our experiments. Instead, we opted for less model complexity, by employing a single, shared linear projection in the multi-head formulation of CAP.
@@
Finally, from Tables 11 and 12 we observe that there are consistent improvements brought by multi-head “pre-aggregation LayerNorm” (cf. Equations (8a), (8b), as the last element in the CAP architecture), compared to no normalization. To check the robustness of CAP to such normalization, we tested the sensitivity of our models to an alternative form of LayerNorm, shown in Appendix E. This alternative is a standard LayerNorm without multi-headness, akin to the post-attention LayerNorm adopted in the transformer model, applied after bag aggregation and after all heads are concatenated in Equations (8a) and (8b). Henceforth, we call it “post-aggregation LayerNorm”. The results in Appendix E show similar performance of CAP-based models when the pre-aggregation LayerNorm is replaced with the post-aggregation LayerNorm, and thus demonstrate the robustness of CAP to different forms of output normalization.
@@ 7 Conclusions and future research
We introduce multi-instance verification (MIV), which combines, but differs from, verification and MIL, and show theoretically and empirically that a lack of information about the query instance in attention is undesirable when pooling the target bag into a bag-level representation. We present a new approach named “Cross Attention Pooling” (CAP) that explicitly considers the query in attention, along with two new attention functions within the CAP framework. We also evaluate a simple baseline model and several benchmark models, adapted from SOTA attention-based MIL methods, that omit the query in attention. Results on three different tasks show that the benchmark methods are not better, and sometimes significantly worse, than the simple baseline. In contrast, the CAP-based methods outperform the baseline and benchmarks by sizeable margins in terms of both classification and key instance detection performance. Ablation studies confirm the superior ability of the new attention functions to identify key instances and establish the contributions of the key components of the CAP architecture.
@@
The abstract representation of “multi-instance verification”, as distinct from traditional MIL, opens up the potential for research on a broad range of problems and applications. For example, either the “query” or the “target” may be a single instance, bag of instances, or bag of bags, and they may exhibit different modalities—for example, one may be textual and the other may consist of images. We believe our work provides a generic framework that may inspire research in directions such as new architectures, attention mechanisms, real-world applications, or methods for selecting key instances and evaluating their quality.
@@
As a first step towards tackling MIV, our paper also offers future opportunities connecting to some domain-specific, well-established areas under active research. For example, while this paper considers CV datasets that enable well-controlled experiments, disentangling factors without compromising on the difficulty of tasks, applications of MIV on more complex datasets such as ImageNet or MS-COCO may lead to research in areas of cross-domain few-shot learning, weakly-supervised one-shot object detection, to name a few. For the NLP domain, a comparison of our approach to other methods specifically designed for particular tasks, e.g., FEVER, may also be an interesting topic for prospective research (cf. Section 2). Beyond CV and NLP, MIV and its solutions may also be applied to potentially broader domains such as time series, multi-modal learning, and so forth. Those generalizations are sufficiently significant and challenging to warrant dedicated, standalone future research, in which domain-specific baselines may be appropriate for comparison, rather than the domain-agnostic MIL models adopted in this paper. Another avenue for future work is to develop methods that scale to large bags or real-time scenarios. Finally, interpretability and its evaluation, in a way that is more general than the “key instance detection” performance considered in this paper, is yet another avenue to be explored. Those research topics, typically investigated in the context of domain-specific applications, are left to future work.
Acknowledgments and Disclosure of Funding
We thank the anonymous reviewers for their helpful comments that enabled us to improve our paper. X. Xu acknowledges financial support from the Machine Learning Group of the Computer Science Department at the University of Waikato.
@@ Appendix A Proof of Proposition 4
Proof [Proposition 4: Variables incorporated in the attention scores]
@@
This concludes the proof of Inequality (4). By symmetry, Inequality (5) can also be shown in a similar manner, which completes the proof of Proposition 4.
@@ Appendix B Details of data and training process
Prior to full training of all network parameters, we adopt an initial learning stage where we freeze the feature extractor’s model weights and train only the parameters specific to the baseline and benchmarks. This initial phase is stopped when the accuracy on the validation dataset does not improve for two epochs.
Subsequently, when all parameters are trained end-to-end, we also adopt early stopping based on validation accuracy, with different stopping criteria for different tasks as discussed below. The optimization for training is conducted using the RMSprop optimizer, with the same parameters for all models: $\rho=0.9,\epsilon=$1e-7. Due to a well-known behavior of “batch normalization” (BatchNorm) causing degradation in performance during transfer learning, see, e.g., Kanavati and Tsuneki (2021), we freeze the BatchNorm layers, if any, in a feature extractor, typically for CV tasks. More precisely, the BatchNorm moving average and the trainable parameters are not updated during training. [^15] We also tried allowing parameters to be trainable but not updating the moving average parameters and found similar experimental results.
For each task, we conducted three rounds of experiments for all models considered in our study. All experiments were run on a cluster of four NVIDIA RTX A6000 GPUs, of which the duration varies depending on early stopping triggers. Roughly speaking, a single round of training of one model takes approximately one hour for the QMNIST task, two to three hours for the FEVER task, and three to five hours for the signature verification task.
All models were developed using Tensorflow 2.9.3, with some use of “TensorFlow official models” 2.9.0 (Yu et al. (2020), Apache License 2.0) and scikit-learn 1.2.0 (Pedregosa et al. (2011), BSD License). The original license and terms-of-use of the assets (i.e., raw data, pre-trained feature extractors) used in our research are listed alongside each asset in the following.
@@ B.1 QMNIST handwriting verification
We collected QMNIST data (BSD-style license, https://github.com/facebookresearch/qmnist/blob/main/LICENSE) according to Yadav and Bottou (2019) and from links therein. QMNIST was created to be similar in structure to MNIST, with a training set composed of 60,000 images of digits (“Train/60K”) and a test set of 10,000 digits (“Test/10K”). These two sets turn out to have non-overlapping writer-IDs, which is ideal for our requirement that test writer’s handwriting is not seen during training. We constructed a training dataset by randomly selecting from the “Train/60K” set, including both the query and target bag. To construct the validation and test datasets, we first split “Test/10K” into two sets with non-overlapping writer-IDs and then drew random samples from each set. By construction, the sample sizes of the train/dev/test datasets are 21,509/2,408/2,253 respectively (the same sample sizes for all rounds of experiments were used to draw random samples), and the proportions of their exemplar labels are approximately 50:50. In an exemplar, the bag size $N$ varies between 3 and 25, with a mean of $6.9$ instances per target bag and a variance of $6.4$ approximately.
We employed ResNet-18 (He et al. (2016)) as the feature extractor, initialized by a set of pre-trained weights based on ImageNet. The pre-trained weights were downloaded from https://pypi.org/project/image-classifiers (MIT License). Prior to inputting the QMNIST images into ResNet-18, we resized them from their original size of $(28\times 28)$ to $(32\times 32)$, because this is the minimum input size required by ResNet-18. After putting an image through ResNet-18, we obtained its feature vector by applying a “Global Average Pooling” (GAP) on the output of the feature extractor’s penultimate layer.
The learning rate of the RMSprop optimizer was piece-wise constant: 1e-4 for the first 5 epochs, 5e-5 for the next 15 epochs, and 2e-5 for the remaining epochs. The mini-batch size was 768, and the early-stopping criterion was non-improvement of validation accuracy for 30 epochs. The number of heads was two for multi-head attention, when applicable.
@@ B.2 Signature verification against multiple anchors
We collected raw images of signatures, both authentic and forged, based on Liwicki et al. (2011) and from the link http://www.iapr-tc11.org/mediawiki/index.php/ICDAR_2011_Signature_Verification_Competition_(SigComp2011) (data disclaimer: http://www.iapr-tc11.org/dataset/ICDAR_SignatureVerification/SigComp2011/disclaimer.pdf). We resized all signature images from their original sizes to $(512\times 256)$ and converted them from RGB to black-and-white. To construct the train, validation, and test datasets for each round of the experiment, we first split the raw data into three sets with non-overlapping writer-IDs and then drew random samples from each set. The sample size for training is around 82,000, and for validation/test is close to 10,000. By construction, the proportion of exemplar labels in all datasets is approximately 50:50. In an exemplar, the bag size $N$ varies between two and eight.
For the feature extractor, we employed one from the EfficientNetV2 family (Tan and Le (2021)), called “EfficientNetV2B3”, initialized with ImageNet-pretrained weights. The pretrained weights were downloaded from https://www.tensorflow.org/versions/r2.9/api_docs/python/tf/keras/applications/efficientnet_v2/EfficientNetV2B3 (CC-BY-4.0 license). We resized all the signature images to an identical size of $(256\times 128)$. After putting an image through EfficientNetV2B3, we obtained its feature vector by applying global average pooling on the output from the penultimate layer.